OpenStudy (anonymous):

Calculate the derivative of the function. Using Chain Rule r(x) = (0.1x^2-4.2x+9.5)^1.5

OpenStudy (kropot72):

\[r \prime(x)=\frac{d(0.1x ^{2}-4.2x+9.5)^{1.5}}{d(0.1x ^{2}-4.2x+9.5)}\times \frac{d(0.1x ^{2}-4.2x+9.5)}{dx}\]

OpenStudy (kropot72):

@chrislb22 Can you find the derivative of each of the terms on the right hand side?

OpenStudy (anonymous):

i think its 0.01-4.2

OpenStudy (kropot72):

I think we must revise some differentiation first: \[y=x ^{1.5}\] \[\frac{dy}{dx}=?\]

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

hey what just happend?

OpenStudy (kropot72):

\[y=x ^{1.5}\] \[\frac{dy}{dx}=1.5x ^{(1.5-1)}=1.5x ^{0.5}=1.5\sqrt{x}\] Do you follow?

OpenStudy (anonymous):

yes i do

OpenStudy (kropot72):

Then you should be able to differentiate the first of the two terms on the right hand side of the equation in my first post above: \[\frac{d(0.1x ^{2}-4.2x+9.5)^{1.5}}{d(0.1x ^{2}-4.2x+9.5)}=?\]

OpenStudy (anonymous):

im sorry im having a hard time figuring out what to do

OpenStudy (anonymous):

Let u = 0.1x² - 4.2x + 9.5 -> u' = ... ( u^n ) ' = n u ^( n -1) * u'

zepdrix (zepdrix):